OpenStudy (anonymous):

What is the simplified form of the expression the square root of the quantity 4 times c to the sixteenth power. − the square root of c squared. + 6the square root of c squared. + the square root of the quantity 9 times c squared.? A:2c4 + 8c B:2c4 + 10c C:2c8 + 8c D:2c8 + 10c

OpenStudy (campbell_st):

\[\sqrt{4c^{16}} - \sqrt{c^2} + 6\sqrt{c^2} + \sqrt{9c^2}\] can you simplify any part of this question...?

OpenStudy (anonymous):

you can change the square root of 4c^16 to 2c^4 ?

OpenStudy (campbell_st):

you can... but there is more to do

OpenStudy (anonymous):

then the the square root of 9 is 3 but i dont know what the exponent would be

OpenStudy (campbell_st):

well all you need to solve is \[\sqrt{c^2} = ?\]

OpenStudy (campbell_st):

do you know about index laws...? \[(c^{16})^{\frac{1}{2}}\]

OpenStudy (campbell_st):

as \[\sqrt{4c^{16}} ... does..\not.. = 2c^4\] its just the power of c that is incorrect.

OpenStudy (campbell_st):

so you have \[2c^4 - \sqrt{c^2} + 6\sqrt{c^2} + 3\sqrt{c^2}\]

OpenStudy (anonymous):

Do i combine like terms ?

OpenStudy (campbell_st):

yes you can... but you need to simplify the square root of c^2

OpenStudy (campbell_st):

here is a something you need to know \[\sqrt{c^2} = c......... \sqrt{4c^{16}} = 2c^8\]

OpenStudy (anonymous):

what was the answer for this question?